How much pure acid should be mixed with 2 gallons of a 50% acid solution in order to get an 80% acid solution?
Answer by checkley75(3666)

sorry for the mistake. I used 5% instead of 50%.
below is the corrected solution & 3 is the answer.
---------------------------------------------------------------------
.5*2+X=.8(X+2)
1+X=.8X+1.6
X-.8X=1.6-1
.2X=.6
X=.6/.2
X=3 GALLONS NEEDED OF THE PURE ACID.
PROOF
.5*3+3=.8(3+2)
1.5+3=.8*5
4=4